Computation of taxable income. The following information relates to Tom, a single taxpayer, age 18. Wages = $7,000 Taxable interest income = 425 Itemized deductions = 310 a. Compute Tom's taxable income assuming he is self-supporting. b. Compute Tom's taxable income assuming he is a dependent of his parents. Tax Year 2019. Please show accurate calculations.
Answer #1

ANSWER

a).

Self supporting : Adjusted Gross Income = 7000 + 425 = 7425

Now, for 2019

Standard deductions = 12200 ; So Max{310,12200} = 12200; so consider standard deductions instead of itemized.

Then , Taxable income = 7425 - 12200 = -4775 < 0 ; since this is negative , implies taxable income is ZERO.

b).

Dependent : Adjusted Gross Income = 7000 + 425 = 7425

Now, for 2019

Standard deductions = AGI + 350 = 7775 ; So Max{7775,1100} = 7775

So consider standard deductions instead of itemized.

Then,

Taxable income = 7425 - 7775 = -350 < 0 ; since this is negative , implies taxable income is ZERO.

__________________________________

If you have any query or any Explanation please ask me in the comment box, i am here to helps you.please give me positive rating.

*****************THANK YOU*************